HUBBLE SPACE TELESCOPE DAILY REPORT #945 PERIOD COVERED: 0800 EST Daily Status Report as of 218/1200Z. 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed FGS 5055 (A Critical Test fo the Galactic Escape Velocity at R(SUN): Cycle 1 OBS) The Fine Guidance Sensors were used to make astrometric observations of the high velocity star G166-37 to determine its galactic escape velocity. The FGSs will be used in the POS mode to measure the target's trigonometric parallax and proper motion. The astrometric observations were monitored in the OSS, and no problems were seen. 1.2 Completed WFC 4292 (Determination of the Masses and Densities of Pluto and Charon Cycle 3) The Wide Field Camera was used to make two observations of the planet Pluto, its moon Charon, and a nearby reference standard star. These images will be used to determine the motion of Pluto about the Pluto/Charon barycenter which will make it possible to compute the mass ratio of the two bodies and put constraints on their masses and densities. Both of the observations were received and displayed. The images were reported to look normal. 1.3 Uplinked Updated Ephemeris Tables, Per MOSES 2871 1.4 Completed FOS 4342 (Damped Lyman Alpha Absorption Lines From Low-Redshift Galaxies - Cycle 3 Medium) The Faint Object Spectrograph was used to obtain a spectra of quasar PKS1451-37 which was culled from the IUE archives to measure its Lyman Alpha absorption features. The observations were executed as scheduled, and no problems were reported. 1.5 Completed WFC 4381 (HST Medium - Deep Survey - Cycle 3) The Wide Field Camera was used to make two set of medium-deep high latitude sky survey observations taken in parallel with the prime science observations. The observations were executed as scheduled, and no problems were reported. 1.6 Completed PC 3872 (Blue Stragglers in the Cores of Globular Clusters) The Planetary Camera was used to make images of the globular cluster NGC-5139 to identify blue straggler stars. The observations was made using two different filters, so that the color index could be found. The OSS monitored the observation and assessed the data quality. No problems were reported. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 3 Successful : 3 Scheduled Reacquisitions : 10 Successful : 10 Losses of Lock : 0 2.2 FHST Updates: Scheduled : 4 Successful : 4 2.3 Significant HSTARs: None 2.4 Operations Note: An attempt to remove the uniform layer of contaminants from the WF/PC CCD windows, which has been accumulating over the last year, was performed by executing a real-time Flash Decontamination procedure. The twenty step procedure was carried out in four uplink windows between 214/1818Z and 215/0026Z. The WF/PC CCD, Hot Junction and Radiator temperatures were monitored throughout (as downlinks permitted) and compared against the anticipated thermal performance to decide which if any pre-planned contingency paths were to be taken (none were) and how the command timing should be spaced within the allowable tolerances. The operations went very much as planned, internal flats were taken to assess how the procedures worked. The decontamination operations appeared to be successful. 3.0 ACTIVITIES SCHEDULED FOR NEXT REPORTING PERIOD: 3.1 Complete Three Sets of WFC 4292 (Determination of the Masses and Densities of Pluto and Charon - Cycle 3) 3.2 Complete Three Sets of HSP/POL 1097 (X-Ray Binaries Part 2 - Cycle 2 High) 3.3 Complete WFC 4848 (GTO Low Latitude Parallel Program - Cycle 3 High) 3.4 Complete PC 4712 (WF/PC Cycle 2 Calibration: Measles Monitor) 3.5 Complete FOS/BL 4976 (UV Rotational Light Curves for Pluto and Charon's UV Spectrum - Split) 3.6 Complete FOS/RD 4123 (FOS Spectral Flat Fields Monitor, Revised) 3.7 Uplink Updated Ephemeris Tables, Per MOSES 2871 3.8 Complete GHRS/2 3845 (Deciphering the UV Emission Lines in the T Tauri systems) 3.9 Complete HSP/UV1 1091 (Ultraviolet Pulsations From X-Ray Pulsars) 3.10 Complete FGS 5055 (A Critical Test of the Galactic Escape Velocity at R(SUN): Cycle 1 Obs.) 4.0 ANOMALY STATUS: 4.1 Significant New Anomalies: None 4.2 Significant Unresolved Anomalies: None 5.0 SIGNIFICANT FORTHCOMING EVENTS: Nothing at this time. /DMH -END-